How do you solve this?
(3/4)x+(1/4)y=(3/2)
-x-(1/3)y=-2
I think the anser is 0=0 True statment, Is that correct?

Answer by RAY100(1637)

To remove fractions multiply 1st eqn by 4 and 2nd by 3
.
(1) 3x +1y =6
.
(2) -3x -1y =-6,,,,,,,,,,now multiply (2) by (-1)
.
(2a) 3x +1y =6,,,,,,,,,,this shows that (1) is the same as (2a),,,or coincident, and the lines overlap,, and there are infinite solutions (infinite coordinates that satisfy either line)
.
.
When we have coincident lines , we do get a true statement, without variables, which means that any variable(x) will work, with the correct corresponding variable (y) of course.
.
Therefore, your 0=0 was correct. and the answer is as stated above.